OpenStudy (anonymous):

A hockey puck slides off the edge of a table with an initial velocity of 26.0 m/s. The height of the table above the ground is 2.00 m. What is the angle below the horizontal of the velocity of the puck just before it hits the ground?

OpenStudy (anonymous):

Can someone please help me with this question...

OpenStudy (darklight):

I think it's 6.87 degrees below the horizontal and just before it hits the ground it should have a downward speed of 6.2524 meters per second

OpenStudy (anonymous):

These are the options a. 12.8° b. 13.5° c. 31.8° d. 77.2° e. 72.6°

OpenStudy (darklight):

Do you have options for the downward speed?

OpenStudy (anonymous):

yes its above the text

OpenStudy (anonymous):

find the final velocity and break it into its components (x,y) the use arctangent for angle

OpenStudy (anonymous):

intial velocity = 26i there is no acceleration in horizontl direction. the initial velocity in vertical direction = 0 after if has fallen let it have a velocity v v^2 - u^2 = 2*a*s v^2 = 2 * 10 * 2 v^2 =40 v = 2 * sqrt 10 so the final velocity has its vertical component of 2 * sqrt 10 since there is no acceleration in horizontal direction the horizontal component is 26 so the angle made with horizontal= tan (theta) = 2*sqrt 10 / 26 theta = tan^-1 ( sqrt 10 / 13)

OpenStudy (anonymous):

is my answer right?